A bank's checkable deposits are $1,340, its loans are $1,100 and the bank has reserves of $240. If the bank faces a required reserve ratio of 13%, then what are the bank's current excess reserves?

a) $122.5

b) $120

c) $65.8

d) $60.5

Bank of the South has checkable deposits of $4,650 and faces a required reserve ratio of 12%. On the asset side of balance sheet the bank has loans and reserves. If the bank is currently holding excess reserves of $122, the the bank's total reserves must be at _____ and the value of its loans must be _____.

a) 680; 3,970

b) 680; 3,740

c) 522; 3,970

d) 522; 3,740
Answer #1

Q1
Answer

required reserve =required reserve ratio * deposits =1340*0.13
=174.2
excess reserves =total reserves - required reserves
=240-174.2
=65.8
the excess reserve is $65.8

Option c
=====
Q2
Answer
required reserve =required reserve ratio * deposits=4650*0.12
=558
total reserves =excess reserves + required reserves=122+558=680
loan =deposits -total reserves
=4650-680
=3970
Option a